OpenStudy (handsomechickenlover):

What is the slope of the line whose equation is −48=2x−8y? WILL MEDAL AND FAN

OpenStudy (handsomechickenlover):

pls help

OpenStudy (samantha2234):

@Handsomechickenlover

OpenStudy (handsomechickenlover):

yeah?

OpenStudy (shansha1):

-48 = 2x - 8y -24 = x - 4y 4y = x + 24 y = (1/4)x + 6 Slope of the line 1/4

OpenStudy (handsomechickenlover):

Thank you sooooooo much and also for the xplaining heres a medal and fan thank you :D